# Beyond Pixels: Efficient Dataset Distillation via Sparse Gaussian Representation

### v1

@@ Abstract
Dataset distillation has emerged as a promising paradigm that synthesizes compact, informative datasets capable of retaining the knowledge of large-scale counterparts, thereby addressing the substantial computational and storage burdens of modern model training. Conventional approaches typically rely on dense pixel-level representations, which introduce redundancy and are difficult to scale up. In this work, we propose GSDD, a novel and efficient sparse representation for dataset distillation based on 2D Gaussians. Instead of representing all pixels equally, GSDD encodes critical discriminative information in a distilled image using only a small number of Gaussian primitives. This sparse representation could improve dataset diversity under the same storage budget, enhancing coverage of difficult samples and boosting distillation performance. To ensure both efficiency and scalability, we adapt CUDA-based splatting operators for parallel inference and training, enabling high-quality rendering with minimal computational and memory overhead. Our method is simple yet effective, broadly applicable to different distillation pipelines, and highly scalable. Experiments show that GSDD achieves state-of-the-art performance on CIFAR-10, CIFAR-100, and ImageNet subsets, while remaining highly efficient encoding and decoding cost. Our code is available at https://github.com/j-cyoung/GSDatasetDistillation.
@@
Guided by the principles of scaling laws Kaplan et al. (2020), deep learning has advanced along the trajectory of larger models, larger datasets, and longer training, which has significantly pushed the performance boundaries of modern models. However, this paradigm also incurs enormous demands on computation, storage, and communication resources, creating a barrier that hinders both broad adoption and sustainable development. Consequently, a pivotal question in the field is how to achieve strong performance under limited computational and data resources.
Against this backdrop, dataset distillation offers a promising solution. Its goal is to distill the essential knowledge of massive datasets into a compact set of synthetic samples, enabling models trained on these small sets to approximate the performance of those trained on the full dataset. By adopting dataset distillation, the costs of model training, data storage, and data transmission can be substantially reduced. Moreover, it opens up a wide range of applications, including efficient data replay for continual learning Yang et al. (2023); Gu et al. (2024), critical data transfer in federated learning Huang et al. (2024a), and privacy preservation Dong et al. (2022).
Dataset distillation optimizes the synthetic data itself, with the objective of minimizing the training-aware information gap between the synthetic and original datasets. Broadly, dataset distillation can be decomposed into two complementary modules: distillation algorithms and data parameterization. The former defines how training-aware key information, such as training gradients Zhao et al. (2021), model parameters Cazenavette et al. (2022), or distributional statistics Zhao & Bilen (2023), is extracted, while the latter determines the representation format of the synthetic data. Although extensive research has focused on innovating distillation algorithms, exploration of the representation space has been relatively limited, with most approaches directly optimizing in the pixel space. Conventional methods typically adopt a naïve pixel grid as the parameterization scheme. Such dense representations fail to capture the relative importance of pixels and lead to substantial storage redundancy. Decoder-based parameterizations, in contrast, can leverage structural priors to enhance image realism and share common information, but they inevitably restrict the optimization space and introduce additional computational and memory overhead. More recently, approaches based on implicit neural representations (INRs), such as DDiF Shin et al. (2025), have substantially reduced per-image storage and thereby improved the diversity of distilled datasets. However, their inherent per-pixel query decoding mechanism incurs prohibitive computational costs, particularly in high-resolution or large-batch scenarios, creating a severe performance bottleneck.
@@ 2.1 Related Work
Modern deep learning models typically rely on large-scale training datasets, whereas the goal of dataset distillation is to replace the dataset with a much smaller set of high-quality training samplesWang et al. (2018). Unlike data selection, dataset distillation explicitly optimizes the synthetic data to achieve superior performance.
@@
The storage footprint of $\mathcal{S}$ is typically required to be far smaller than that of $\mathcal{T}$. In dataset distillation, the synthetic dataset is treated as a set of learnable objects, where the pixel values of distilled images $s_{j}$ are regarded as optimization parameters. The ultimate goal is to train a model $\theta_{\mathcal{S}}$ on the compact distilled dataset such that its performance closely matches that of a model $\theta_{\mathcal{T}}$ trained on the full dataset Wang et al. (2018):
@@
This optimization objective can be approached via meta-learning, but such methods often consume substantial GPU memory Feng et al. (2024) and yield suboptimal performance due to weak supervision signals. To address these challenges, many alternative distillation objectives have been proposed. The core idea is to extract and align critical training-aware information between $\mathcal{S}$ and $\mathcal{T}$. Mainstream approaches include trajectory matching(TM) Cazenavette et al. (2022); Du et al. (2023); Liu et al. (2025); Guo et al. (2024); Cui et al. (2023), which constrains the distance between model parameters trained on the two datasets; distribution matching(DM) Zhao & Bilen (2023); Zhao et al. (2023); Liu et al. (2023); Wei et al. (2024); Wang et al. (2025b), which aligns feature distributions; and gradient matching(DC) Zhao et al. (2021); Lee et al. (2022); DU et al. (2023), which constrains the gradients derived from the two datasets. Despite the drastic reduction in storage requirements, distilled datasets often achieve competitive training performance, and the resulting synthetic images can diverge significantly from the originals. Consequently, dataset distillation has found applications in continual learning Yang et al. (2023); Gu et al. (2024), privacy preservation Dong et al. (2022); Chung et al. (2024); Zheng et al. (2025), federated learning Huang et al. (2024a); Jia et al. (2025); Yan et al. (2025), and neural architecture search Such et al. (2020).
@@
The parameterization of the distilled dataset $\mathcal{S}$ critically influences both performance and efficiency. Early methods directly optimized RGB pixels, but this dense representation wastes storage by ignoring spatial redundancy. To address this, decoder-based approaches Liu et al. (2022); Liu & Wang (2023); Cazenavette et al. (2023) learn latent codes with a shared decoder, which reduces redundancy but limits flexibility in optimization and adds computation costs. Dictionary-based methods Deng & Russakovsky (2022); Wei et al. (2023) instead combine basis images with index matrices to support domain adaptation and generalization, though they still store pixel-level bases and are restricted by linear combinations.
Recent studies have introduced compact low-level representations to reduce image redundancy and increase distilled samples within a fixed storage budget, such as downscaling Kim et al. (2022), frequency-domain transforms Shin et al. (2023), and implicit neural representations (INRs) Shin et al. (2025). INR-based methods show strong performance by encoding each image as a neural field with fewer parameters, but their efficiency is limited since decoding requires per-pixel RGB queries, causing heavy overhead for high-resolution and large-batch training.
@@
Gaussian splatting Kerbl et al. (2023) represents 3D scenes using Gaussian ellipsoids, offering higher rendering efficiency and explicit geometric control compared to NeRF-based representations Mildenhall et al. (2021). Originally developed for novel view synthesis, it initializes a set of Gaussian primitives within a 3D scene, which are then rendered from multiple viewpoints and optimized to match the corresponding ground-truth images, ultimately enabling arbitrary-view synthesis. In recent years, Gaussian splatting has been extended to a wide range of domains, including digital humans Qian et al. (2024); Zielonka et al. (2025), autonomous driving Zhou et al. (2024); Huang et al. (2024b), dynamic scene modeling Wu et al. (2024); Huang et al. (2024c), and 3D editing Chen et al. (2024). More recently, Gaussian representations have also been applied to 2D vision tasks such as image super-resolution Hu et al. (2024); Chen et al. (2025), image representation Zhu et al. (2025); Weiss & Bradley (2024); Zhang et al. (2025), and video representation Wang et al. (2025a), where Gaussian primitives are used to efficiently encode high-fideliy image of videos with fine-grained textures.
@@ 2.2 Motivation
Gaussian representations enable faster decoding and rendering, especially compared to the state-of-the-art DDiF Shin et al. (2025) method. DDiF relies on implicit neural representations using a SIREN network Sitzmann et al. (2020) to model distilled images, where each pixel’s RGB value must be queried individually, as illustrated in Figure 1a. This approach incurs substantial time and memory overhead as the number and resolution of distilled images increase. In contrast, GSDD implements a custom CUDA operator for batched rendering of Gaussians. This allows all Gaussians to be efficiently rendered into distilled images in a single pass. As shown in Figure 1b, this design significantly reduces both inference and optimization time and memory usage, thereby improving the scalability of dataset distillation methods.
@@ 3.1 Parameterization of Dataset Distillation By Gaussian Splatting
Thus, each Gaussian component can be fully described by a 9-dimensional vector $p_{k}=(u_{k},v_{k},l_{11}^{k},l_{21}^{k},l_{22}^{k},\mathbf{c}_{k},\alpha_{k})\in\mathbb{R}^{9}.$ Accordingly, the parameter set of an entire distilled image $s_{j}$ is given by $\mathbf{p}^{j}=\{p_{k}^{j}\}_{k=1}^{M}$, and the parameterization of the whole distilled dataset $\mathcal{S}$ is $\mathcal{P}_{\mathcal{S}}=\{\mathbf{p}^{j}\}_{j=1}^{N_{\mathcal{S}}}.$ Based on this parameterization, the rendering process of a distilled image can be expressed as a differentiable synthesis function
@@ 3.2 Efficient Gaussian Splatting Design
Existing open-source rasterization operators are primarily designed for 3D scene reconstruction or single-image tasks such as representation learning and super-resolution, and generally lack support for multi-image rendering. To fully leverage the parallelism of modern GPUs, we implement customized data structures and rendering kernels for both the forward and backward passes. Specifically, all Gaussian primitives across the entire distilled dataset are represented as a single contiguous 1D vector, and a globally unique ID (GUID) system is used to assign threads that rasterize multiple images in parallel. Details are provided in Appendix A.3, and code will be open-sourced.
@@ 3.3 Analysis of the Gaussian Representation
Table 1: Results on ImageNet-Subset, TM
@@
Table 2: Results with DC/DM
@@ 4.1 Experimental Results
We evaluate our method on standard datasets: CIFAR-10, CIFAR-100 (at $32\times 32$ resolution), and six ImageNet subsets (at $128\times 128$ resolution). The general pipeline of dataset distillation involves two stages: (1) generating a synthetic dataset, and (2) training a model from scratch on the synthetic data. The performance of the trained model on the real test set is used to assess the quality of the distilled dataset. We include a wide range of recent dataset distillation baselines, including TM Cazenavette et al. (2022), FRePo Zhou et al. (2022), IDC Kim et al. (2022), FreD Shin et al. (2023), HaBa Liu et al. (2022), RTP Deng & Russakovsky (2022), HMN Liu et al. (2022), SPEED Wei et al. (2023), NSD Yang et al. (2025), GLaD Cazenavette et al. (2023), H-GLaD Zhong et al. (2025), LD3M Moser et al. (2024) and DDiF Shin et al. (2025).
@@
We use TM Cazenavette et al. (2022), DM Zhao & Bilen (2023), and DC Zhao et al. (2021) as the underlying distillation algorithms in our experiments. All parameters of Gaussians are trained using the Adam optimizer, and the learning rate for Gaussian parameters is uniformly set to 0.001. For CIFAR datasets, we apply ZCA whitening as a standard preprocessing step. We fix the boundary loss weight to $\lambda_{\text{boundary}}=0.1$. Experiments are conducted on NVIDIA V100 and RTX 4090 GPUs. To initialize the Gaussian representation, we randomly sample real images from the original dataset and fit them with Gaussians before entering the standard distillation loop. We provide complete implementation details and hyperparameters in the Appendix A.1.
Table 3: Ablation

### v2

@@ Abstract
Dataset distillation has emerged as a promising paradigm that synthesizes compact, informative datasets capable of retaining the knowledge of large-scale counterparts, thereby addressing the substantial computational and storage burdens of modern model training. Conventional approaches typically rely on dense pixel-level representations, which introduce redundancy and are difficult to scale up. In this work, we propose GSDD, a novel and efficient sparse representation for dataset distillation based on 2D Gaussians. Instead of representing all pixels equally, GSDD encodes critical discriminative information in a distilled image using only a small number of Gaussian primitives. This sparse representation could improve dataset diversity under the same storage budget, enhancing coverage of difficult samples and boosting distillation performance. To ensure both efficiency and scalability, we adapt CUDA-based splatting operators for parallel inference and training, enabling high-quality rendering with minimal computational and memory overhead. Our method is simple yet effective, broadly applicable to different distillation pipelines, and highly scalable. Experiments show that GSDD achieves state-of-the-art performance on CIFAR-10, CIFAR-100, and ImageNet subsets, while remaining highly efficient encoding and decoding cost. The code is available at https://github.com/j-cyoung/GSDatasetDistillation.
@@
Guided by the principles of scaling laws kaplan2020scalinglawsneurallanguage, deep learning has advanced along the trajectory of larger models, larger datasets, and longer training, which has significantly pushed the performance boundaries of modern models. However, this paradigm also incurs enormous demands on computation, storage, and communication resources, creating a barrier that hinders both broad adoption and sustainable development. Consequently, a pivotal question in the field is how to achieve strong performance under limited computational and data resources.
Against this backdrop, dataset distillation offers a promising solution. Its goal is to distill the essential knowledge of massive datasets into a compact set of synthetic samples, enabling models trained on these small sets to approximate the performance of those trained on the full dataset. By adopting dataset distillation, the costs of model training, data storage, and data transmission can be substantially reduced. Moreover, it opens up a wide range of applications, including efficient data replay for continual learning yang2023an; guSummarizingStreamData2024, critical data transfer in federated learning huang2024overcoming, and privacy preservation dongPrivacyFreeHow2022.
Dataset distillation optimizes the synthetic data itself, with the objective of minimizing the training-aware information gap between the synthetic and original datasets. Broadly, dataset distillation can be decomposed into two complementary modules: distillation algorithms and data parameterization. The former defines how training-aware key information, such as training gradients zhaoDatasetCondensationGradient2021, model parameters Cazenavette_2022_CVPR, or distributional statistics zhaoDatasetCondensationDistribution2023, is extracted, while the latter determines the representation format of the synthetic data. Although extensive research has focused on innovating distillation algorithms, exploration of the representation space has been relatively limited, with most approaches directly optimizing in the pixel space. Conventional methods typically adopt a naïve pixel grid as the parameterization scheme. Such dense representations fail to capture the relative importance of pixels and lead to substantial storage redundancy. Decoder-based parameterizations, in contrast, can leverage structural priors to enhance image realism and share common information, but they inevitably restrict the optimization space and introduce additional computational and memory overhead. More recently, approaches based on implicit neural representations (INRs), such as DDiF shin2025distilling, have substantially reduced per-image storage and thereby improved the diversity of distilled datasets. However, their inherent per-pixel query decoding mechanism incurs prohibitive computational costs, particularly in high-resolution or large-batch scenarios, creating a severe performance bottleneck.
@@ 2.1 Related Work
Modern deep learning models typically rely on large-scale training datasets, whereas the goal of dataset distillation is to replace the dataset with a much smaller set of high-quality training sampleswangDatasetDistillation2018. Unlike data selection, dataset distillation explicitly optimizes the synthetic data to achieve superior performance.
@@
The storage footprint of $\mathcal{S}$ is typically required to be far smaller than that of $\mathcal{T}$. In dataset distillation, the synthetic dataset is treated as a set of learnable objects, where the pixel values of distilled images $s_{j}$ are regarded as optimization parameters. The ultimate goal is to train a model $\theta_{\mathcal{S}}$ on the compact distilled dataset such that its performance closely matches that of a model $\theta_{\mathcal{T}}$ trained on the full dataset wangDatasetDistillation2018:
@@
This optimization objective can be approached via meta-learning, but such methods often consume substantial GPU memory fengEmbarrassinglySimpleDataset2024 and yield suboptimal performance due to weak supervision signals. To address these challenges, many alternative distillation objectives have been proposed. The core idea is to extract and align critical training-aware information between $\mathcal{S}$ and $\mathcal{T}$. Mainstream approaches include trajectory matching(TM) Cazenavette_2022_CVPR; duMinimizingAccumulatedTrajectory2023; liuDatasetDistillationAutomatic2025; guoLosslessDatasetDistillation2024; cuiScalingDatasetDistillation2023, which constrains the distance between model parameters trained on the two datasets; distribution matching(DM) zhaoDatasetCondensationDistribution2023; zhaoImprovedDistributionMatching2023; liuDREAMEfficientDataset2023a; Wei_2024_CVPR; Wang_2025_CVPR, which aligns feature distributions; and gradient matching(DC) zhaoDatasetCondensationGradient2021; leeDatasetCondensationContrastive2022; duuentialSubsetMatching2023, which constrains the gradients derived from the two datasets. Despite the drastic reduction in storage requirements, distilled datasets often achieve competitive training performance, and the resulting synthetic images can diverge significantly from the originals. Consequently, dataset distillation has found applications in continual learning yang2023an; guSummarizingStreamData2024, privacy preservation dongPrivacyFreeHow2022; chung2024rethinking; zheng2025improvingnoiseefficiencyprivacypreserving, federated learning huang2024overcoming; UPFL; yan2025fedvck, and neural architecture search such2020generative.
@@
The parameterization of the distilled dataset $\mathcal{S}$ critically influences both performance and efficiency. Early methods directly optimized RGB pixels, but this dense representation wastes storage by ignoring spatial redundancy. To address this, decoder-based approaches liuDatasetDistillationFactorization2022; liuMGDDMetaGenerator2023; cazenavetteGeneralizingDatasetDistillation2023 learn latent codes with a shared decoder, which reduces redundancy but limits flexibility in optimization and adds computation costs. Dictionary-based methods dengRememberDistillingDatasets2022; weiSparseParameterizationEpitomic2023 instead combine basis images with index matrices to support domain adaptation and generalization, though they still store pixel-level bases and are restricted by linear combinations.
Recent studies have introduced compact low-level representations to reduce image redundancy and increase distilled samples within a fixed storage budget, such as downscaling kimDatasetCondensationEfficient2022, frequency-domain transforms shinFrequencyDomainBasedDataset2023, and implicit neural representations (INRs) shin2025distilling. INR-based methods show strong performance by encoding each image as a neural field with fewer parameters, but their efficiency is limited since decoding requires per-pixel RGB queries, causing heavy overhead for high-resolution and large-batch training.
@@
Gaussian splatting 3dgs represents 3D scenes using Gaussian ellipsoids, offering higher rendering efficiency and explicit geometric control compared to NeRF-based representations NeRF. Originally developed for novel view synthesis, it initializes a set of Gaussian primitives within a 3D scene, which are then rendered from multiple viewpoints and optimized to match the corresponding ground-truth images, ultimately enabling arbitrary-view synthesis. In recent years, Gaussian splatting has been extended to a wide range of domains, including digital humans qian20243dgs; zielonka2025drivable, autonomous driving zhou2024drivinggaussian; huang2024textit, dynamic scene modeling wu20244d; huang2024sc, and 3D editing chen2024gaussianeditor. More recently, Gaussian representations have also been applied to 2D vision tasks such as image super-resolution hu2024gaussiansrhighfidelity2d; chen2025generalizedefficient2dgaussian, image representation zhu2025largeimagesgaussianshighquality; weiss2024gaussianbillboardsexpressive2d; gaussianimage, and video representation wang2025gsvcefficientvideorepresentation, where Gaussian primitives are used to efficiently encode high-fidelity image of videos with fine-grained textures.
@@ 2.2 Motivation
Gaussian representations enable faster decoding and rendering, especially compared to the state-of-the-art DDiF shin2025distilling method. DDiF relies on implicit neural representations using a SIREN network NEURIPS2020_53c04118 to model distilled images, where each pixel’s RGB value must be queried individually, as illustrated in Figure 1a. This approach incurs substantial time and memory overhead as the number and resolution of distilled images increase. In contrast, GSDD implements a custom CUDA operator for batched rendering of Gaussians. This allows all Gaussians to be efficiently rendered into distilled images in a single pass. As shown in Figure 1b, this design significantly reduces both inference and optimization time and memory usage, thereby improving the scalability of dataset distillation methods.
@@ 3.1 Parameterization of Dataset Distillation By Gaussian Splatting
Thus, each Gaussian component can be fully described by a 9-dimensional vector $p_{k}=(u_{k},v_{k},l_{11}^{k},l_{21}^{k},l_{22}^{k},\mathbf{c}_{k},\alpha_{k})\in\mathbb{R}^{9}.$ Accordingly, the parameter set of an entire distilled image $s_{j}$ is given by $\mathbf{p}^{j}=\{p_{k}^{j}\}_{k=1}^{M}$, and the parameterization of the whole distilled dataset $\mathcal{S}$ is $\mathcal{P}_{\mathcal{S}}=\{\mathbf{p}^{j}\}_{j=1}^{N_{\mathcal{S}}}.$ Based on this parameterization, $R$ can be seen as a rendering function, which maps the parameters $\mathbf{p}^{j}$ onto a predefined pixel coordinate grid $\mathcal{C}$, and is implemented as the cuda operator,
@@ 3.2 Efficient Gaussian Splatting Design
Existing open-source rasterization operators are primarily designed for 3D scene reconstruction or single-image tasks such as representation learning and super-resolution, and generally lack support for multi-image rendering. To fully leverage the parallelism of modern GPUs, we implement customized data structures and rendering kernels for both the forward and backward passes. Specifically, all Gaussian primitives across the entire distilled dataset are represented as a single contiguous 1D vector, and a globally unique ID (GUID) system is used to assign threads that rasterize multiple images in parallel. Details are provided in Appendix C, and code will be open-sourced.
@@ 3.3 Analysis of the Gaussian Representation
Table 1: Test accuracy on ImageNet subsets ($128\times 128$) with trajectory matching (TM).
@@
Table 2: Test accuracy on ImageNet subsets ($128\times 128$) with gradient matching (DC) and distribution matching (DM).
@@ 4.1 Experimental Results
We evaluate our method on standard datasets: CIFAR-10, CIFAR-100 (at $32\times 32$ resolution), six ImageNet subsets (at $128\times 128$ resolution and $256\times 256$ resolution), and ImageNet-1K (at $224\times 224$ resolution). The general pipeline of dataset distillation involves two stages: (1) generating a synthetic dataset, and (2) training a model from scratch on the synthetic data. The performance of the trained model on the real test set is used to assess the quality of the distilled dataset. We include a wide range of recent dataset distillation baselines, including TM Cazenavette_2022_CVPR, FRePo zhouDatasetDistillationUsing2022, IDC kimDatasetCondensationEfficient2022, FreD shinFrequencyDomainBasedDataset2023, HaBa liuDatasetDistillationFactorization2022, RTP dengRememberDistillingDatasets2022, HMN liuDatasetDistillationFactorization2022, SPEED weiSparseParameterizationEpitomic2023, NSD NSD, GLaD cazenavetteGeneralizingDatasetDistillation2023, H-GLaD Zhong_2025_CVPR, LD3M moser2024latentdatasetdistillationdiffusion and DDiF shin2025distilling.
@@
We use TM Cazenavette_2022_CVPR, DM zhaoDatasetCondensationDistribution2023, and DC zhaoDatasetCondensationGradient2021 as the underlying distillation algorithms in our experiments. All parameters of Gaussians are trained using the Adam optimizer, and the learning rate for Gaussian parameters is uniformly set to 0.001. For CIFAR datasets, we apply ZCA whitening as a standard preprocessing step. We fix the boundary loss weight to $\lambda_{\text{boundary}}=0.1$. Experiments are conducted on NVIDIA V100 and RTX 4090 GPUs. To initialize the Gaussian representation, we randomly sample real images from the original dataset and fit them with Gaussians before entering the standard distillation loop. We provide complete implementation details and hyperparameters in the Appendix A.
Storage Budget and Evaluation Protocol
We evaluate performance under varying storage budgets, specified in IPC (images per class). Since our method utilizes multiple low-storage Gaussian images, we additionally report GPC (Gaussian Images per Class) in Appendix A. The number of Gaussians per image is computed as $\text{pts}=\frac{\text{res}\times\text{res}\times 3\times\text{IPC}\times 2}{\text{GPC}\times 9}$, where the factor of 2 accounts for the use of bf16, and the denominator 9 reflects the number of parameters per Gaussian.
Performance Comparison
Through the integration of GSDD, our method consistently outperforms the pixel-based TM baseline and other dataset parameterization methods on six ImageNet subsets with IPC=1, as shown in Table 2. The results on the higher-resolution ImageNet subset (256×256) are provided in Table 4, where our method remains superior at higher resolutions. In addition, we also report experiments conducted at a lower resolution on CIFAR-10/100, with the results presented in Appendix Table 13. When compared to the state-of-the-art method DDiF, our approach achieves consistent improvements, demonstrating highly competitive performance. These results highlight the effectiveness of GSDD in enhancing the quality of distilled datasets. We further report results on ImageNet-1K($224\times 224$) in Table 17 in Appendix B.4.
Universality to Matching Objectives
Cross-Architecture Performance
An essential property of dataset distillation is robustness across model architectures. The distilled data should retain high performance even when the evaluation model differs from the one used during distillation. To assess this, we synthesize data using a ConvNet and evaluate it on a variety of downstream architectures, including ResNet, VGG, AlexNet and ViT. The results, summarized in Table 4 and Table 14, demonstrate that the increased diversity introduced by GSDD leads to substantial improvements in cross-architecture generalization.
Table 3: Results on ImageNet-subsets with Distribution Matching (DM) for $256\times 256$
Method Nette Woof Fruit Yellow Meow Squawk
Vanilla 32.1 20.0 19.5 33.4 21.2 27.6
IDC 53.7 30.2 33.1 52.2 34.6 47.0
FreD 54.2 31.2 32.5 49.1 34.0 43.1
LatentDD 56.1 28.0 30.7 36.3 47.1
DDiF 67.8 39.6 43.2 63.1 44.8 67.0
GSDD 70.0 42.6 51.2 67.4 46.4 70.4
Table 4: Cross-Architecture Performance
Subset Nette Woof Fruit Yellow Meow Squawk
TM 22.0 14.8 17.1 22.3 16.2 25.5
IDC 27.9 19.5 23.9 28.0 19.8 29.9
FreD 36.2 23.7 23.6 31.2 19.1 37.4
GLaD 30.4 17.1 21.1 19.6 28.2
H-GLaD 30.8 17.4 21.5 20.1 28.8
LD3M 32.0 19.9 21.4 22.1 30.4
DDiF 59.3 34.1 39.3 51.1 33.8 54.0
GSDD 58.1 34.6 39.9 53.6 34.0 58.0
4.2 Benchmark Rendering Efficiency
To evaluate the efficiency and representational capacity of GSDD, we conduct an image-fitting experiment comparing several recent parameterization methods with the same storage budget, including FreD shinFrequencyDomainBasedDataset2023, SPEED weiSparseParameterizationEpitomic2023 and DDiF shin2025distilling. Specifically, we use the ImageNette dataset and fit the first 100 images from each class (1,000 images in total) at a resolution of 128. The optimization is performed for 1,000 steps, and for each method we perform a grid search over learning rates and optimizers to ensure a fair comparison. As shown in Table 5, GSDD shows great performance across representational quality (PSNR), GPU memory usage, and runtime efficiency. High fidelity and low computational overhead of GSDD enables fast synthesis of more diverse (Higher GPC) distilled datasets, which in turn leads to improved downstream performance. We further verify in Section 4.3 that larger GPC values with fixed IPC generally correspond to stronger distilled data performance.
Table 5: Comparison of Parameterization Methods on Image Fitting Efficiency and Representational Quality
FreD SPEED DDiF GSDD
PSNR(dB)↑ 15.69 17.77 23.23 23.22
VRAM(MB)↓ 280.4 298.5 1078.8 150.2
Time(s)↓ 15.5 69.4 1088.4 24.5
We further conduct a comprehensive quantitative comparison with the state-of-the-art DDiF method in terms of both decoding and training performance. As shown in Figure 5, we measure the inference time and memory consumption of both methods across varying per-image storage budgets, image resolutions, and batch sizes. The results demonstrate that GSDD consistently outperforms DDiF in both inference/training speed and memory efficiency. This advantage becomes especially pronounced when handling high-resolution inputs or large batch sizes, where GSDD achieves several-fold, or even an order-of-magnitude, reductions in both computation time and memory usage. These findings indicate that our method not only maintains strong image representation capabilities for distillation but also offers superior rendering speed and significantly lower memory footprint, thereby substantially improving the scalability of dataset distillation methods in practical deployment scenarios.
4.3 Ablation Study
Table 6: Ablation Study on Individual Components